OpenStudy (anonymous):

Boy A throws a ball vertically up with speed u . Boy b on the top catches it and throws down with u to A . the times for the two travels (ascent : decent) are in ratio 2:1 The height of the tower is ?

OpenStudy (anonymous):

I think it's quite easy question, you must be able to solve that. Try to do this yourself.

OpenStudy (anonymous):

is that a maximum height...

OpenStudy (anonymous):

@ArchiePhysics

OpenStudy (anonymous):

First you need to find the time of ascent using v =vi-gt

OpenStudy (anonymous):

i got u=10t

OpenStudy (anonymous):

@ArchiePhysics .............

OpenStudy (anonymous):

h=12 u^2/25g is this correct answer?

OpenStudy (anonymous):

@open_study1

OpenStudy (anonymous):

yes i have an option like that

OpenStudy (anonymous):

but how u got it???

OpenStudy (ajprincess):

vat r ur other options?

OpenStudy (anonymous):

I have used standard formulas h(t) = hi+vit+at^2/2 v = vi+at and given ratio t1/t2=2

OpenStudy (anonymous):

Try to solve it yourself. It is easy.

OpenStudy (anonymous):

6u^2/25g 12u^2/25g 18u^2/25g 24u^2/25g

OpenStudy (anonymous):

@ajprincess these are my options

OpenStudy (anonymous):

i got u=10t @ArchiePhysics is this correct

OpenStudy (anonymous):

so ......wat is the real answer??

OpenStudy (anonymous):

@ArchiePhysics

OpenStudy (anonymous):

can u show ur method @ajprincess

OpenStudy (ajprincess):

 used these three eqns and found it.